How many gallons of maple syrup at $5 a gallon and how many gallons of corn syrup at 70 cents a gallon must be mixed to make 100 gallons to sell at $1.99 a gallon?
Answer by lwsshak3(11628)

How many gallons of maple syrup at $5 a gallon and how many gallons of corn syrup at 70 cents a gallon must be mixed to make 100 gallons to sell at $1.99 a gallon?
**
let x=amt of maple syrup to be mixed
100-x=amt of corn syrup to be mixed
..
5x+.70(100-x)=1.99*100
5x+70-.70x=199
4.3x=129
x=30
100-x=70
amt of maple syrup to be mixed=30 gal
amt of corn syrup to be mixed=70 gal